OpenStudy (kise_ryouta):

a=11+4b−4c solve for c.

alones (alones):

Alright so flip the equation. \(\ 4b−4c+11=a\)

alones (alones):

Alright so add \(\ -11\) to both sides. \[\ 4b−4c+11+−11=a+−11\] \[\ 4b−4c=a−11\] So than add \(\ -4b\) to each side \[\ 4b−4c+−4b=a−11+−4b\] \[\ −4c=a−4b−11\]

alones (alones):

Your last step woule be dividing both sides by \(\ -4.\) \[c=\frac{ -1 }{ 4 }a=\frac{ a−4b−11 }{ -4 }\]

alones (alones):

So your final answer should be \[c=\frac{ -1 }{ 4 }a+b+\frac{ 11 }{ 4 }\]